# The Role of Privacy Guarantees in Voluntary Donation of Private Data for Altruistic Goals

### v1

@@ 3 Methodology
As mentioned in section 1, we address research questions RQ1 and RQ2 through a user survey. This section details our survey’s design (subsection 3.1) and the cognitive interview process used to verify respondents’ comprehension of the presented scenarios (subsection 3.2). We also describe our participant selection strategy (subsection 3.3) and the limitations of our methodology (subsection 3.5).
3.1 Survey Structure
Our survey design is shown in Figure 1.
@@ 3.1 Survey Structure
Figure 1: Survey design. Each participant is presented with a donation scenario with a survey introduction with two variations: recipient entity and disease. statements (PG(1)–PG(4) and AG(1)–AG(2)) are also presented in the donation scenario. Participants are then asked five sets of questions (survey questions). When the control statements are presented, the trust (dashed) is omitted. The willingness to donate and trust questions (dark shaded) are associated with an open-text question. The scenario understanding is used as a filter. Participants who did not understand the donation scenario were excluded from the analysis.
@@
First, we present to the respondents a suvery introduction, which we detail in Figure 2.
@@
Figure 2: Survey Introduction. The entity type (for-profit or non-profit) and the disease (selected from a list of common chronic diseases) are randomly selected.
We present a donation scenario using this introduction: we want to assess respondents’ willingness to donate their health data to a recipient entity developing a new treatment for a specific disease. The type of entity (for-profit or non-profit) and the disease (cancer, diabetes, heart failure, high blood pressure, and stroke) are randomly selected; the disease options are taken from a list of common chronic diseases published by the Centers for Medicare and Medicaid Services (CMS) [36].
Then, we present to the respondents either a control (no statement) or an experimental statement. Experimental statements are composed of either one privacy statement, alone, or one privacy statement and one auditing statement. Privacy statements are uniformly selected at random from a pool of four privacy statements; auditing statements are selected at random from two auditing statements. We have written these statements to align with PG(1)–PG(4) (see section 1), and we list them in Figure 3. We further distinguish the auditing guarantee into two different auditing statements to align with AG(1)–AG(2), and we list them in Figure 4.
@@
Figure 3: Privacy statements PG(1)–PG(4). One of the five statements (including the control statement) is randomly presented in the donation scenario immediately after the survey introduction.
@@
Figure 4: Auditing statements AG(1)–AG(2). One of the three statements (including the control statement) is randomly presented in the donation scenario after the privacy statement.
@@
  • demographics and experiences, including age, gender, education level, donation history, and technical background.
@@
Figure 5: To measure privacy expectations respondents reported their agreement with each statement listed above on a 4-point Likert Scale from “Very Likely” to “Very Unlikely”.
@@
We also ask respondents to explain their responses to the willingness to donate and trust questions.These explanations are collected through open-ended questions: “Why your are willing (or unwilling) to share your medical record with this entity?” and “Please explain why you do (or do not) trust that the entity will handle your data as described.” (We note that the scenario understanding and trust questions are only asked when a non-control privacy statement is presented).
3.2 Questionnaire Refinement
@@
The final questionnaire was refined through multiple iterations. Before launching the survey, we prescreened 49 users on the Prolific online platform [1] by asking them to complete a short demographics questionnaire and indicate their willingness to participate in a cognitive interview. Then, we conducted 17 cognitive interviews to ensure participants understood the questionnaire correctly. Each interview involved presenting the questionnaire to participants and actively seeking feedback on their interpretation of the data donation scenario, and the privacy and auditing statements. We additionally presented all statements to the interviewees and asked for their understanding and perception of each of them. We continued refining the questionnaire through cognitive interviews until no further constructive feedback was received. In the final interview, we observe a clear understanding of the statements and the donation scenario. For example, the participant described anonymization (PG(1)) as "eliminating any sort of demographic or personal information affiliated with your data that could be associated with you if the data were to somehow be leaked" and access control (PG(2)) as "some sort of database that only certain individuals have access to by using a passcode and certain credentials to to log in." Each Prolific user who completed a cognitive interview was paid \$15.
3.3 Survey Sampling
Following best practice in security research [102], we recruited 560 respondents via Prolific. We restricted the respondents to adults residing in the U.S. to obtain focused results. Then, we ensured a balanced gender distribution: 49.11% men, 48.57% women, 1.6% non-binary, and 0.7% unwilling to disclose. Each participant received $0.83\$$ upon completing the questionnaire; the median completion time was around 7 minutes.
@@ 3.3 Survey Sampling
Out of these 560 respondents, we sequentially excluded the following from analysis: 13 that reported non-binary gender [^2] As we lack statistical power to draw meaningful conclusions about the privacy sentiments of the non-binary community specifically, we do not include these data in our analysis. However, this in no way diminishes the importance of understanding the unique and valuable insights of this population, which should be explored with either larger samples (yielding statistical power) or targeted research specific to their viewpoints and perspectives [39, 38].; 31 that submitted incomplete responses; 20 that failed our attention check (subsection 3.1); and 11 that indicated they did not understand the survey scenario (using scenario understand). Thus, our final dataset comprises 485 respondents, whose demographics are reported in Table I. How many respondents were assigned to each condition is reported in Table II.
@@
Age $\geq$ 40 219 45.15%
< 40 166 54.85%
Gender Woman 254 52.37%
Man 231 47.63%
Education w/ BSc degree 352 72.58%
w/o BSc degree 133 27.42%
Technical Background Yes 128 26.39%
No 357 73.61%
Donation History Yes 55 11.34%
No 430 88.66%
TABLE I: Participant demographics. We note that the second category of each demographical attribute (e.g., no technical background) is considered the baseline scenario during analysis.
@@
PG(1) 22 18 19 19 19 18
PG(2) 17 18 19 18 18 19
PG(3) 18 20 18 18 17 17
PG(4) 21 18 18 20 18 19
Ctrl. - - 19 - - 20
TABLE II: Number of respondents assigned to each condition (i.e., who saw each privacy statement and auditing statement, or no privacy statement (bottom row) or no auditing statement (columns 4 and 6).
3.4 Analysis
We analyzed the open-text questions about willingness to donate and trust (see subsection 3.1) using inductive-thematic open coding. Two researchers independently coded each entry and generated a codebook from a random sample of at least 100 (20.6%) responses. Then, they composed a final codebook and double-coded all responses. Since all responses were double-coded and inconsistencies were resolved, we do not report inter-rater reliability (IRR) [69].
For all quantitative analyses, we binarized the responses to the willingness to donate, privacy expectations, and trust questions (see subsection 3.1)) to get ‘True’ if the respondents selected ‘Likely’ or ‘Very Likely’, and ‘False’ otherwise.
@@ 3.4 Analysis
For RQ2, the dependent variable was willingness to donate, and the independent variables were the presence of a given privacy statement, the privacy expectation for each guarantee, as well as demographics and experiences. We categorized education into two groups: with and without a bachelor’s degree. We also categorized age into two groups and cut to the age of forty. We took “no bachelor’s degree” and “less than forty years old” as the reference categories, respectively. For gender, we took man as the reference category. Technical background and donation history are binary factors, and we took the negative response as the reference category. We built separate regression models for the two recipient entities: for-profit and non-profit.
3.5 Limitations
Although the four privacy-preserving guarantees examined in this study (section 1) span many PETs, they are not fully comprehensive and do not capture the full complexity of PETs nor of real-world threats, which may involve privacy issues beyond the scope of our investigation. Moreover, despite our efforts to mitigate misunderstandings through rigorous cognitive interviews, filtering out respondents who reported not fully understanding the scenario, and controlling for privacy expectations in our statistical analyses, the brief descriptions of each guarantee might have led to partial or incorrect understanding among respondents, potentially influencing our results. While this study employs vignette scenarios that can effectively mirror real-world behaviors [90], discrepancy between participants’ expressed privacy concerns and their actual behaviors may exist (i.e., privacy paradox [13]).
@@ 3.5 Limitations
Furthermore, statistically-significant results do not inherently guarantee the absence of underlying biases that could skew the findings. In fact, various unmeasured factors that were not controlled for could influence the interpretations of the results. Finally, our respondents were recruited via Prolific, and thus, may not accurately reflect the diversity of the entire U.S. population, let alone the varied perspectives of individuals from other countries. This limitation challenges the generalizability of our findings.
4 Results
4.1 RQ1: Understanding of PGs
@@
Next, we answer the first research question (RQ1, see section 1): “How well do people understand what is offered by the privacy-preserving guarantees PG(1)–PG(4) and auditing guarantees AG(1)–AG(2)?". We analyze and report the survey responses and evaluate the rationales for the responses.
@@
Analysis Set Up.Analysis Set UpAnalysis Set Up.
We compare privacy expectations on PG(1)—PG(4) and AG(1)–AG(2) among groups of respondents who received different privacy statements (listed in Figure 3) and auditing statements (listed in Figure 4). In Table III, we show the privacy expectations of respondents who were shown different privacy statements.
To assess the statistical significance of the descriptive quantitative results presented above, we used binomial logistic regression to analyze the relationship between the presence of a privacy statement in the scenario and respondents’ privacy expectations. We summarize results, which we distinguish for non-profit and for-profit entities, in Table IV.
Relationship between Privacy Statement Shown and Privacy Expectation Statement Shown Overall For-Profit Non-Profit PG(1) Anonymization 0.68 0.68 0.67 + AG(1) Expert audit 0.63 0.59 0.68 + AG(2) Self audit 0.65 0.67 0.63 Control 0.38 0.26 0.50 PG(2) Access control 0.50 0.58 0.42 + AG(1) Expert audit 0.51 0.35 0.67 + AG(2) Self audit 0.50 0.56 0.44 Control 0.49 0.37 0.60 PG(3) Data expiration 0.83 0.67 1.00 + AG(1) Expert audit 0.92 0.83 1.00 + AG(2) Self audit 0.86 0.85 0.88 Control 0.41 0.37 0.45 PG(4) Purpose restriction 0.78 0.72 0.84 + AG(1) Expert audit 0.71 0.62 0.80 + AG(2) Self audit 0.81 0.67 0.94 Control 0.56 0.37 0.75 AG(1) Expert audit 0.82 0.78 0.87 Control 0.62 0.58 0.66 AG(2) Self audit 0.53 0.50 0.57 Control 0.26 0.25 0.28
TABLE III: Percentage of respondents who had a positive privacy expectation when shown a particular privacy statement in their donation scenario. The values in the table are the percentage of respondents in a given condition who had the privacy expectation (see Figure 5) that corresponded to the privacy statement they were presented. The 2nd column (“overall") reports results across both entities (for-profit and non-profit), the 3rd column reports the results of the subset in the for-profit entity conditions, and the 4th column reports the results of the subset in the non-profit entity conditions. (Recall that agreement has been binarized as discussed in subsection 3.4). For example, the left-most and top-most numerical cell indicates that 68% of participants in both entity scenarios who were shown PG(1) alone – with no auditing statement – expected that their data would be anonymized.
Statement Shown OR 95% CI $p$-value
For-Profit PG(1) Anonymization 5.07 [1.69, 17.52] 0.006
PG(2) Access control 1.71 [0.6, 5.23] 0.325
PG(3) Data expiration 6.29 [2.09, 20.44] 0.001
PG(4) Purpose restriction 3.43 [1.19, 10.59] 0.026
AG(1) Expert audit 2.59 [1.33, 5.20] 0.006
AG(2) Self audit 3.04 [1.59, 5.93] <0.001
Non-Profit PG(1) Anonymization 1.95 [0.69, 5.57] 0.208
PG(2) Access control 1.27 [0.45, 3.61] 0.651
PG(3) Data expiration 16.7 [3.64, 120.49] <0.001
PG(4) Purpose restriction 2.04 [0.55, 7.11] 0.266
AG(1) Expert audit 3.41 [1.59, 7.85] 0.002
AG(2) Self audit 3.41 [1.79, 6.61] <0.001
TABLE IV: Effect of presenting different privacy statements on respondents’ privacy expectations toward the two recipient entities. OR: odd ratio; CI: confidence interval.
@@
Response Distribution and Statistical Analysis Results.Response Distribution and Statistical Analysis ResultsResponse Distribution and Statistical Analysis Results.
We find that privacy expectations differ based on the data-collecting entity. In the control group, which was not shown any privacy statement, 26%-37% of respondents expected for-profit entities to provide anonymization, data expiration, access control, use restriction to a specific purpose, expert auditing, and self-auditing. In contrast, a higher percentage of respondents (45%-75%) expected non-profit entities to employ these mechanisms. For instance, we note that respondent P430 (who received a for-profit scenario) expected privacy protections even though they were in the control group (and thus received no statement about privacy protection): "I also assume that the data is looked at in the aggregate and likely no one at the company knows me.". In the same situation, respondent P53 (who received a non-profit scenario instead) made a stronger assumption: "I trust that the organization will uphold strict privacy and ethical standards".
In the for-profit scenario, all privacy statements, except for the access-control statement, significantly increase the corresponding privacy expectations ($p$-value < $0.05$). In the non-profit scenario, where expectations are already high, only the privacy statement about data expiration significantly increases expectations (from 60% to 100%). Comparing the for-profit and non-profit subsets of the descriptive results (2nd and 3rd column of Table III), we observe that presenting a privacy statement was particularly beneficial in for-profit scenarios. Indeed, the presence of a privacy statement raised respondents’ privacy expectations to nearly the same level as non-profit scenarios for anonymization (68% and 67%) and access control (58% and 42%). It also reduced the gap in expectations for data expiration (67% and 100%)) and purpose restriction (72% and 84%)).
In both scenarios, the auditing statements increase the expectation that a given privacy guarantee will be audited ($p$-value < $0.05$). However, we note that incorporating an audit statement does not markedly change privacy expectations. The statistical significance of the auditing statements in Table IV indicates that respondents shown the audit statement correctly expected the corresponding type of audit, while respondents that were not shown the auditing statement did not expect this protection. As observed in  Table III, presenting an auditing statement does not increase the number of respondents who expect the privacy guarantee implied by the privacy statement they were shown, despite the stronger assurances that audits imply.
@@
Qualitative Analysis of Privacy Expectations via Open-Answer Responses.Qualitative Analysis of Privacy Expectations via Open-Answer ResponsesQualitative Analysis of Privacy Expectations via Open-Answer Responses.
Satisfaction with the privacy guarantees. 16.3% and 20.8% of respondents in the for-profit and non-profit scenarios, respectively, were convinced by the received privacy statement. Some respondents, like P56, vaguely stated that the policy looks promising: "I see their policy, and they have to follow their own policy.". Others felt explicitly safer and trusted the statement. For example, P494 noted: "Trust is very important when it comes to medical data. I believe the organization has privacy policies that outline that they collect and will use my data. I also believe the organization will employ security measures to safeguard data". Similarly, P555 highlighted the presence of the privacy statement: "(…) purposely states there is software in place to conserve privacy." and P516 felt their information was safe because of it: "(…) with the privacy protection in place, they are isolating the data they need while basically "throwing out" the rest by putting in under that protection. In essence, my information is safe, and they’re only using what they said they’d use". P208 also felt safer contributing to research knowing the privacy mechanisms were in place: "I want to be able to contribute to research to better improve cancer treatment, and I feel safe if my data is protected through the mechanisms above".
Belief in legal and reputational obligations. 16.3% and 10.6% of respondents in the for-profit and non-profit scenarios, respectively, believed that entities are forced to protect the donated data due to legislative requirements and reputational concerns. For example, P28 stated regarding for-profits: "A for-profit organization wouldn’t want to violate HIPAA, HITECH laws". Additionally, P117 wrote "I would trust them to do the right thing so they won’t face lawsuits". Regarding non-profits, P141 noted: "I trust that the law will restrict any data leaks to third parties".
Trust in the recipient entity. 13.9% and 17.5% of respondents in the for-profit and non-profit scenarios, respectively, expressed a general trust in the recipient entity without specifying reasons. For instance, P70 succinctly said "I feel they are reliable and trustworthy" and P50 stated that "I assume they take their research seriously, so they would handle the data carefully". Other respondents, like P145, reported having no reason not to trust it: "I have no reason to think they would do anything nefarious with my medical data".
Trust in the auditing process. 7.8% and 7.9% of respondents in the for-profit and non-profit scenarios, respectively, found the auditing statement to add at least some reliability. For example, P202 showed some reservation but found confidence in the public nature of audits: "I don’t fully trust them, but I somewhat do, particularly if audits and verification of results are made public. That said, claiming that the advisory board is external is only partly reassuring, as it’s appointed by the institute." Additionally, P540 had trust in the entity’s data handling because "they let outsiders audit them" and P51 (who received the expert-auditing statement) because of "safelocks and checks in place". On the other hand, one interviewee doubted the expert auditing process and stated: "somebody else says something doesn’t mean that it’s real.". No questionnaire respondent explicitly raised concern towards the auditing process.

### v2

@@ 3. Methodology
As mentioned in § 1, we address research questions RQ1 and RQ2 through a user survey. In this section, we discuss the ethical considerations of our study, outline the survey design (§ 3.2), and describe the cognitive interview process used to ensure participants’ understanding of the presented scenarios (§ 3.3). We also describe our participant selection strategy (§ 3.4) and the limitations of our methodology (§ 3.6).
3.1. Ethics
Before we began our research, our ethics review board approved all of the study methods, including the cognitive interviews and the survey. A consent form detailing the data storage, handling, and survey compensation is presented at the beginning of the survey. Participants can opt out of disclosing any demographic data without influencing their compensation in the survey.
Each participant received $1.20\$$ upon completing the survey. To determine this figure, we ran a test of 20 participants, who took 5m56s (median) to complete it. Thus, the actual compensation was $12\$$/hr, which aligns with Prolific recommendations. For the signup survey of the cognitive interview, we set the study name to “Sign Up for Paid Interview on Data Donation” with a payment of $1.25\$$ ($13.68\$$/hr based on a loose completion estimate of 5 minutes). However, most respondents refused our interview request. Thus, in order to encourage participants interested in attending the interview to complete the screener, we lowered the compensation for the screening survey to $9.96\$$/hr. The interview reward is $15\$$, and all interviews lasted at most thirty minutes.
The survey is conducted through Qualtrics (Qualtrics, 2020), with the IP address collection feature disabled to protect respondents’ privacy. Interviews are conducted via Zoom (zoom, 2024), and participants are instructed to avoid using their real names when joining the meeting to ensure anonymity. Research participants are anonymously recruited through Prolific (Prolific, 2024), as described in greater detail below.
3.2. Survey Structure
@@
Figure 1. Survey design. Each participant is presented with a donation scenario and a survey introduction with two variations: recipient entity and disease. statements (PG(1)–PG(4) and AG(1)–AG(2)) are also presented in the donation scenario. Participants are then asked five sets of questions (survey questions). When the control statements are presented, the trust (dashed) is omitted. The willingness to donate and trust questions (dark-shaded) are associated with an open-text question. The scenario understanding is used as a filter. Participants who did not understand the donation scenario were excluded from the analysis. A sample screen that participants would encounter is provided in Figure 6.
@@
Our survey design is shown in Figure 1, and the full survey can be found in Appendix A. First, we present to the respondents a suvery introduction, which we detail in Figure 2.
@@
Figure 2. Survey Introduction. The entity type (for-profit or non-profit) and the disease (selected from a list of common chronic diseases) are randomly selected.
We present a donation scenario using this introduction: we want to assess respondents’ willingness to donate their health data to a recipient entity developing a new treatment for a specific disease. The type of entity (for-profit or non-profit) and the disease (cancer, diabetes, heart failure, high blood pressure, and stroke) are randomly selected; the disease options are taken from a list of common chronic diseases published by the Centers for Medicare and Medicaid Services (CMS) (for Medicare and Services, 2024).
Then, we present to the respondents either a control (no statement) or an experimental statement. Experimental statements are composed of either one privacy statement, alone, or one privacy statement and one auditing statement. Privacy statements are uniformly selected at random from a pool of four privacy statements; auditing statements are selected at random from two auditing statements. We have written these statements to align with PG(1)–PG(4) (see § 1), and we list them in Figure 3. We further distinguish the auditing guarantee into two different auditing statements to align with AG(1)–AG(2), and we list them in Figure 4.
@@
Figure 3. Privacy statements PG(1)–PG(4). One of the five statements (including the control statement) is randomly presented in the donation scenario immediately after the survey introduction.
@@
Figure 4. Auditing statements AG(1)–AG(2). One of the three statements (including the control statement) is randomly presented in the donation scenario after the privacy statement.
@@
  • demographics and experiences, including age, gender, education level, donation history, technical background (i.e., education background and job field relationships with computer technologies), and egocentricity(i.e., whether respondents or their close relatives have the disease, as defined in (Grgić-Hlača et al., 2020; Plane et al., 2017)).
@@
Figure 5. To measure privacy expectations respondents reported their agreement with each statement listed above on a 4-point Likert Scale from “Very Likely” to “Very Unlikely”.
@@
We also ask respondents to explain their responses to the willingness to donate and trust questions. These explanations are collected through open-ended questions: “Why you are willing (or unwilling) to share your medical record with this entity?” and “Please explain why you do (or do not) trust that the entity will handle your data as described.” We note that the scenario understanding and trust questions are only asked when a non-control privacy statement is presented.
A sample screen that participants would encounter is provided in Figure 6.
Imagine that a non-profit organization (e.g., university, non-profit research institute) wants to develop a new treatment for Hypertension (High blood pressure). They need medical data from people with and without Hypertension (High blood pressure) to develop the treatment. They ask you to donate your medical record to help develop the treatment. Your medical record contains your: (i) personal information, which may include information about your age, weight, gender, race; (ii) medical history, which may include information about allergies, illnesses, surgeries, immunizations, and results of physical exams and tests; and (iii) medical behavior, which may include information about medicines taken and health habits, such as smoking habits, diet and exercise.
The non-profit organization uses a privacy-preserving technology to ensure that your data is deleted after a set period of time. You will be able to choose from a list of time period options how long your data will be stored. Anyone interested, including you and experts you trust, will be able to verify that the privacy-preserving technology is working as described. Anyone, including you and experts you trust, can post their verification results publically.
How would you rate your understanding of the above scenario? Fully understand Mostly understand Partially understand Not understand
Figure 6. Sample survey screen that composed of three segments. The first segment presents the survey introduction, which varies the disease and the entity among all participants. The second segment includes the randomized privacy statement and auditing statement, with color coding applied to emphasize these elements. The third segment contains the survey question that participants are required to answer.
3.3. Survey Refinement
@@
The final survey was refined through multiple iterations. Before launching the survey, we prescreened 49 users on the Prolific online platform (Prolific, 2024) by asking them to complete a short demographics survey and indicate their willingness to participate in a cognitive interview. Then, we conducted 17 cognitive interviews to ensure participants understood the survey correctly. Each interview involved presenting the survey to participants and actively seeking feedback on their interpretation of the data donation scenario, and the privacy and auditing statements. We additionally presented all statements to the interviewees and asked for their understanding and perception of each of them. We continued refining the survey through cognitive interviews until no further constructive feedback was received. In the final interview, we observe a clear understanding of the statements and the donation scenario. For example, the participant described anonymization (PG(1)) as ”eliminating any sort of demographic or personal information affiliated with your data that could be associated with you if the data were to somehow be leaked” and access control (PG(2)) as ”some sort of database that only certain individuals have access to by using a passcode and certain credentials to to log in.” Our final survey can be found in Appendix A.
3.4. Survey Sampling
Following the recommendation of (Tang et al., 2022), we recruited participants through Prolific, gathering responses from 560 individuals. To ensure focused results, we restricted respondents to adults residing in the U.S. and requested a gender-balanced distribution. Our recruitment included 272 men, 275 women, 9 nonbinary individuals, and 4 participants who chose not to disclose their gender.
@@ 3.4. Survey Sampling
Of these 560 respondents, we sequentially excluded the following from the analysis: 35 that submitted incomplete responses; 20 that failed our attention check (§ 3.2); and 11 that indicated they did not understand the survey scenario (using scenario understand). Thus, our final dataset comprises 494 respondents. To enhance statistical power, we limit the levels of age and education to binary groups based on the data distribution among our participants. We report their demographics in Table 1 and the number of respondents assigned to each condition in Table 2.
@@
Age 18 - 29 120 24.3%
30 - 49 222 44.9%
50 - 64 111 22.5%
65+ 41 8.3%
Gender Woman 254 51.42%
Man 231 46.76%
Others 9 1.82%
Education B.S. or above 358 72.47%
Up to H.S. 136 27.53%
Technical Background Yes 129 26.11%
No 365 73.89%
Donation History Yes 56 11.34%
No 438 88.66%
Egocentricity Yes 201 40.69%
No 293 59.31%
Table 1. Participant demographics. We note that the second category of each demographical attribute (e.g., no technical background) is considered the baseline scenario during analysis, except for “age”, which is treated numerically.
@@
PG(1) 22 18 19 19 20 18
PG(2) 17 21 19 18 18 19
PG(3) 18 20 18 19 19 17
PG(4) 21 18 19 20 18 19
Ctrl. - - 19 - - 21
Table 2. Number of respondents assigned to each condition (i.e., who saw each privacy statement and auditing statement, or no privacy statement (bottom row) or no auditing statement (columns 4 and 7).
3.5. Analysis
We analyzed the open-text questions about willingness to donate and trust (see § 3.2) using inductive-thematic open coding. Two researchers independently coded each entry and generated a codebook from a random sample of at least 100 (20.2%) responses. Then, they composed a final codebook and double-coded all responses. Since all responses were double-coded and inconsistencies were resolved, we do not report inter-rater reliability (IRR) (McDonald et al., 2019).
@@ 3.5. Analysis
For RQ2, the dependent variable was willingness to donate, and the independent variables were the presence of a given privacy statement, the privacy expectation for each guarantee, as well as demographics and experiences. We collapsed the independent variables into a binary measure, designating responses of “Likely” or “Very Likely” as True and all others as False. We did so to avoid the ambiguity introduced by intermediate scale points and to simplify the statistical model.
@@
We categorized education into two groups: with and without a bachelor’s degree. We took “no bachelor’s degree” and “less than forty years old” as the reference categories, respectively. For gender, we took man as the reference category. Technical background and donation history are binary factors, and we took the negative response as the reference category. We built separate regression models for the two recipient entities: for-profit and non-profit.
3.6. Limitations
Although the four privacy-preserving guarantees examined in this study (§ 1) cover a broad range of PETs, they are not fully comprehensive and do not capture the full complexity of PETs or real-world threats, which may involve privacy issues beyond the scope of our investigation. The randomly presented disease may also influence the results. Furthermore, our approach may not fully reflect the intricacies of real-world donation scenarios, where multiple guarantees often coexist. While such complexities lie outside the primary focus of our research, qualitative investigations have been conducted in prior work (Aitken et al., 2016; Mori, 2016), and quantitative exploration of these aspects remains an interesting direction for future research.
Despite our efforts to mitigate misunderstandings—through rigorous cognitive interviews, filtering respondents who reported not fully understanding the scenarios, and controlling for privacy expectations in our statistical analyses—the brief descriptions of each guarantee might still have led to partial or incorrect understanding among respondents. This limitation may have influenced our results.
Additionally, while our study employs vignette scenarios that effectively mirror real-world behaviors (Redmiles et al., 2017), there may still be discrepancies between participants’ expressed privacy concerns and their actual behaviors, a phenomenon commonly referred to as the privacy paradox (Barth and de Jong, 2017).
@@ 3.6. Limitations
Furthermore, statistically significant results do not inherently eliminate the risk of underlying biases that could skew the findings. Various unmeasured factors not accounted for in our analysis might have influenced the interpretations of the results. Lastly, our respondents were recruited via Prolific, which may limit the generalizability of our findings. The sample might not fully represent the diversity of the U.S. population, nor does it capture the perspectives of individuals from other countries, highlighting a key limitation of this work.
4. Results
4.1. RQ1: Understanding of PGs
@@
Next, we answer the first research question (RQ1, see § 1): “How well do people understand what is offered by the privacy-preserving guarantees PG(1)–PG(4) and auditing guarantees AG(1)–AG(2)?”. We analyze and report the survey responses and evaluate the rationales for the responses.
4.1.1. Analysis Setup
We compare privacy expectations on PG(1)—PG(4) and AG(1)–AG(2) among groups of respondents who received different privacy statements (listed in Figure 3) and auditing statements (listed in Figure 4). In Figure 7, we show the privacy expectations of respondents who were shown different privacy statements. To assess the statistical significance of the descriptive quantitative results presented above, we used binomial logistic regression to analyze the relationship between the presence of a privacy statement in the scenario and respondents’ privacy expectations. We summarize results, which we distinguish for non-profit and for-profit entities, in Figure 8.
Figure 7. Percentage of respondents who had a positive privacy expectation when shown a particular privacy statement in their donation scenario. The values in the table are the percentage of respondents in a given condition who had the privacy expectation (see Figure 5) that corresponded to the privacy statement they were presented. The 3rd column (“overall”) reports results across both entities. For example, the right-most and top-most numerical cell indicates that 68% of participants in both entity scenarios who were shown PG(1) alone – with no auditing statement – expected their data to be anonymized. The agreements are binarized here to present the overall tendency.
Figure 8. Odd ratios and 95% confidence intervals of presenting different privacy statements on the respondents’ privacy expectations toward the two kinds of recipient entities, for-profit and non-profit.
4.1.2. Response Distribution and Statistical Analysis Results
We find that privacy expectations differ based on the data-collecting entity. In the control group, which was not shown any privacy statement, 26%-37% of respondents expected for-profit entities to provide anonymization, data expiration, access control, use restriction to a specific purpose, expert auditing, and self-auditing. In contrast, a higher percentage of respondents (43%-76%) expected non-profit entities to employ these mechanisms, except for self-auditing (AG2). For instance, we note that respondent P430 (who received a for-profit scenario) expected privacy protections even though they were in the control group (and thus received no statement about privacy protection): ”I also assume that the data is looked at in the aggregate and likely no one at the company knows me”. In the same situation, respondent P53 (who received a non-profit scenario instead) made a stronger assumption: ”I trust that the organization will uphold strict privacy and ethical standards”.
In the for-profit scenario, all privacy statements, except for the access-control statement (PG2), significantly increase the corresponding privacy expectations ($p$-value ¡ $0.05$). In the non-profit scenario, where expectations are already high, only the privacy statement about data expiration (PG3) significantly increases expectations (from 57% to 100%). Comparing the for-profit and non-profit subsets of the descriptive results (1st and 2nd column of Figure 7), we observe that presenting a privacy statement was particularly beneficial in for-profit scenarios. Indeed, the presence of a privacy statement raised respondents’ privacy expectations to nearly the same level as non-profit scenarios for anonymization (PG1, 68% and 67%) and access control (PG2, 58% and 42%). It also reduced the gap in expectations for data expiration (PG3, 67% and 100%)) and purpose restriction (PG4, 74% and 84%)).
In both scenarios, the auditing statements increase the expectation that a given privacy guarantee will be audited ($p$-value ¡ $0.05$). However, we note that incorporating an audit statement does not markedly change privacy expectations. The statistical significance of the auditing statements in Figure 8 indicates that respondents shown the audit statement correctly expected the corresponding type of audit, while respondents that were not shown the auditing statement did not expect this protection. As observed in  Figure 7, presenting an auditing statement does not increase the number of respondents who expect the privacy guarantee implied by the privacy statement they were shown, despite the stronger assurances that audits imply.
4.1.3. Qualitative Analysis of Privacy Expectations via Open-Answer Responses
@@
Satisfaction with the privacy guarantees..Satisfaction with the privacy guarantees.Satisfaction with the privacy guarantees.. 16.3% and 20.8% of respondents in the for-profit and non-profit scenarios, respectively, were convinced by the received privacy statement. Some respondents, like P56, vaguely stated that the policy looks promising: ”I see their policy, and they have to follow their own policy”. Others felt explicitly safer and trusted the statement. For example, P494 noted: ”Trust is very important when it comes to medical data. I believe the organization has privacy policies that outline that they collect and will use my data. I also believe the organization will employ security measures to safeguard data”. Similarly, P555 highlighted the presence of the privacy statement: ”(…) purposely states there is software in place to conserve privacy”. and P516 felt their information was safe because of it: ”(…) with the privacy protection in place, they are isolating the data they need while basically ”throwing out” the rest by putting in under that protection. In essence, my information is safe, and they’re only using what they said they’d use”. P208 also felt safer contributing to research knowing the privacy mechanisms were in place: ”I want to be able to contribute to research to better improve cancer treatment, and I feel safe if my data is protected through the mechanisms above”.
@@
Belief in legal and reputational obligations..Belief in legal and reputational obligations.Belief in legal and reputational obligations.. 16.3% and 10.6% of respondents in the for-profit and non-profit scenarios, respectively, believed that entities are forced to protect the donated data due to legislative requirements and reputational concerns. For example, P28 stated regarding for-profits: ”A for-profit organization wouldn’t want to violate HIPAA, HITECH laws”. Additionally, P117 wrote ”I would trust them to do the right thing so they won’t face lawsuits”. Regarding non-profits, P141 noted: ”I trust that the law will restrict any data leaks to third parties”.
@@
Trust in the recipient entity..Trust in the recipient entity.Trust in the recipient entity.. 13.9% and 17.5% of respondents in the for-profit and non-profit scenarios, respectively, expressed a general trust in the recipient entity without specifying reasons. For instance, P70 succinctly said ”I feel they are reliable and trustworthy” and P50 stated that ”I assume they take their research seriously, so they would handle the data carefully”. Other respondents, like P145, reported having no reason not to trust it: ”I have no reason to think they would do anything nefarious with my medical data”.
[…]
Trust in the auditing process..Trust in the auditing process.Trust in the auditing process.. 7.8% and 7.9% of respondents in the for-profit and non-profit scenarios, respectively, found the auditing statement to add at least some reliability. For example, P202 showed some reservation but found confidence in the public nature of audits: ”I don’t fully trust them, but I somewhat do, particularly if audits and verification of results are made public. That said, claiming that the advisory board is external is only partly reassuring, as it’s appointed by the institute”. Additionally, P540 had trust in the entity’s data handling because ”they let outsiders audit them” and P51 (who received the expert-auditing statement) because of ”safelocks and checks in place”. On the other hand, one interviewee doubted the expert auditing process and stated: ”somebody else says something doesn’t mean that it’s real”. No survey respondent explicitly raised concern towards the auditing process.